OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -8, and 2 + 5i

OpenStudy (anonymous):

If 2 + 5 i is a root, them 2-5i is also.

OpenStudy (anonymous):

The answer should be \[ (x-4) (x-(2+5 i)) (x-(2-5 i)) (x+8) \] Do you know how to expand it?

OpenStudy (anonymous):

\[ (x-4) (x-(2+5 i)) (x-(2-5 i)) (x+8)=x^4-19 x^2+244 x-928 \]

OpenStudy (anonymous):

what i dont understand is how there is a constant when 5i is part of it

OpenStudy (anonymous):

\[ (x-(2+5 i)) (x-(2-5 i))=x^2-4 x+29 \]

OpenStudy (anonymous):

Just expand it